7x + 3y = -2
-7x - 7y = 14

Answer by mananth(16949)

7 x + 3 y = -2 .............1
-7 x -7 y = 14 .............2
Eliminate y
multiply (1)by 7
Multiply (2) by 3
49 x 21 y = -14
-21 x -21 y = 42
Add the two equations
28 x = 28
/ 28
x = 1
plug value of x in (1)
7 x + 3 y = -2
7 + 3 y = -2
3 y = -2 -7
3 y = -9
y = -3

m.ananth@hotmail.ca